### v1

# Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression

Abstract
Large language models (LLMs) excel across many natural language tasks, yet their generalisation to structural perturbations in logical contexts remains poorly understood. We introduce a controlled evaluation framework that probes reasoning reliability through four targeted stress tests: (1) rule deletion, removing either redundant or essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites generated through several families of equivalence laws (contrapositive, double negation, implication, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that introduces 2–5 simultaneous logical transformations.
Across three representative model families—BERT, Qwen2, and LLaMA-like models—our experiments reveal a strikingly consistent pattern: all models achieve perfect accuracy on the base tasks and remain fully generalise to redundant rule deletion and all equivalence-based rewrites (single or multi-law), but fail sharply under essential rule deletion (dropping to 25% accuracy) and collapse completely in the presence of explicit contradictions (0% accuracy). These results demonstrate that LLMs possess stable invariance to semantic-preserving logical transformations, yet remain fundamentally brittle to missing or conflicting evidence. Our framework provides a clean diagnostic tool for isolating such reasoning failure modes and highlights persistent gaps in the logical generalisation abilities of current LLMs.
Codehttps://github.com/14H034160212/lemo
Introduction
Reasoning over structured rule systems is a fundamental capability required for reliable decision-making and language understanding. Although large language models (LLMs) achieve impressive performance across numerous natural language benchmarks, it is still unclear to what extent these models perform true logical reasoning rather than relying on memorized patterns or superficial correlations. Much prior work evaluates LLMs in settings where all relevant information is present in canonical form; however, real-world reasoning often involves incomplete information, rephrased or transformed rules, redundant evidence, and even explicit contradictions. Understanding how LLMs behave under such perturbations is essential for both theoretical insight and safe deployment.
To address this gap, we introduce a highly controlled evaluation framework that isolates core aspects of rule-based reasoning. The framework is built around a canonical multi-step logical chain (e.g., green → cold → rough → young → nice) that requires models to integrate several rules in sequence. From this base structure, we generate a family of targeted variants that probe different dimensions of reasoning generalisation:
Rule deletion. We remove either redundant rules—whose absence should not change conclusions—or essential rules that constitute the backbone of the inference chain. This tests whether models distinguish between logically important and unimportant components.
Contradictory evidence injection. We introduce explicit facts that conflict with conclusions implied by the rule chain, requiring the model to reconcile or prioritize between mutually inconsistent pieces of information.
Logic-preserving transformations. We rewrite rules using several major families of logical equivalence laws, including contraposition, double negation, implication-to-disjunction, De Morgan’s laws, identity, and commutativity of disjunction. These variants preserve the underlying semantics while changing surface form, allowing us to test whether LLMs track logical content rather than specific linguistic templates.
Multi-law equivalence stacking. We introduce conditions in which 2–5 logical equivalence laws are applied simultaneously to create highly redundant, cluttered, and structurally complex rule sets, all of which remain logically equivalent to the base rules.
We evaluate this framework across three distinct model families—BERT, Qwen2, and LLaMA-like models—using identical training and testing pipelines. Despite architectural differences, all models exhibit the same qualitative and quantitative behavior:
  • Perfect reasoning on the base tasks (100% accuracy)
  • Full generalisation to redundant rule deletion (100% accuracy)
  • Catastrophic failure under essential rule deletion (25% accuracy)
  • Complete collapse under explicit contradictions (0% accuracy)
  • Full invariance to all six classes of logical equivalence rewrites (100% accuracy)
  • Full invariance even under multi-law equivalence stacking (100% accuracy)
Together, these results reveal a consistent pattern: LLMs correctly identify which rules are structurally necessary, generalise flawlessly across a wide spectrum of logically equivalent transformations, but lack the capacity to adjust reasoning when evidence is incomplete or contradictory. Rather than flexibly integrating conflicting information, models tend to complete the rule chain whenever possible, even when doing so contradicts stated facts.
Our contributions provide a clean and extensible diagnostic tool for probing reasoning behavior, enabling precise identification of reasoning failure modes. The findings highlight a persistent gap between surface-level generalisation to paraphrase-like variation and broader generalisation to structural perturbation or conflict, underscoring the need for models and training strategies that can reason reliably under minimal, missing, or inconsistent evidence.
Related Work
Research on logical reasoning in large language models (LLMs) spans multi-step deduction, logic-driven augmentation, abductive explanation, abstract reasoning assessment, generalisation analysis, and hybrid neuro-symbolic methods. A growing body of work shows that LLMs often fail even on elementary logical transformations. For example, the Reversal Curse demonstrates that models trained on “A is B” frequently fail to infer the logically equivalent “B is A,” revealing structural weaknesses in bidirectional generalisation (berglund2023reversal). Such findings motivate deeper investigation into how LLMs internalise logical relationships and under which conditions reasoning fidelity deteriorates.
Early studies like Transformers as Soft Reasoners demonstrated that transformers can perform multi-step inference over natural language rules and generalise to deeper reasoning depths (transformerClark). Subsequent work showed that this generalisation is fragile: Bao et al. found that models trained on short reasoning chains fail to extrapolate to longer or structurally shifted chains, highlighting persistent deficiencies in compositional generalisation (bao2022multi). Similar concerns are echoed in broader analyses showing that LLMs lack strong inductive biases and struggle with systematic, relational, or abstract reasoning (ijcai2024p693; Cheng2025Em; cheng2026sapp).
A parallel line of work attempts to strengthen reasoning via logic-driven augmentation. Wang et al. propose a symbolic context-extension and augmentation framework that applies equivalence laws to derive new training instances, improving model performance on ReClor and LogiQA (wang-etal-2022-logic). Bao et al. extend this idea by using Abstract Meaning Representation (AMR) to generate semantically grounded logical variants that enhance generalisation to unseen reasoning depths (bao2024). Despite their effectiveness, these methods primarily expand data distributions rather than isolating the minimal perturbations that cause reasoning failures.
Beyond deduction, abductive reasoning has been explored through frameworks such as AbductionRules, which train models to generate rules explaining surprising inputs (young-etal-2022-abductionrules). Other studies examine LLM generalisation under structured perturbations: Bao et al. show that reordering, compressing, or restructuring premises leads to substantial degradation in reasoning fidelity (baoAccess). Hybrid neuro-symbolic approaches, such as ChatLogic, integrate LLMs with Prolog-style inference mechanisms to address these limitations, although they rely heavily on external symbolic scaffolds (10650138). Complementing these evaluations, Bao’s dissertation provides a unified perspective on out-of-distribution generalisation, structural generalisation, and equivalence invariance by analysing architecture-agnostic failure modes across premise deletion, contradiction insertion, and logic-preserving rewrites (bao2025thesis).
Across these threads, prior work demonstrates meaningful progress but still lacks frameworks capable of systematically disentangling generalisation to redundant versus essential evidence, contradiction sensitivity, and invariance to different classes of logical equivalence. These unresolved limitations motivate our controlled, fine-grained evaluation framework that isolates minimal structural perturbations and reveals predictable failure modes in multi-step reasoning across LLM architectures.
Methodology
1. Synthetic Logical Environment
Base facts + rule chain generation
$\Downarrow$
2. Structural Variants
V1: Redundant rule removal
V2: Essential rule removal V3: Contradictory facts V4: Logical equivalence laws (1–5)
$\Downarrow$
3. Model Fine-Tuning & Inference
BERT / Qwen2 / TinyLlama (LoRA)
$\Downarrow$
4. Evaluation & Analysis
Accuracy, $\Delta_{\text{base}}$, variant breakdown
Figure 1: Compact system architecture of our logical reasoning generalisation framework. Each base example is transformed into structural variations (rule deletion, contradiction injection, and logic-equivalence rewrites). Models are trained on base data only and evaluated across all variants to measure generalisation and sensitivity.
The system architecture in Figure 1 outlines the full pipeline used to construct, transform, and evaluate logical reasoning tasks. The process begins with the generation of synthetic logical environments containing base facts and a multi-step rule chain designed to support deterministic deductive inferences. From each base environment, we systematically derive multiple structural variants that probe different dimensions of reasoning generalisation: (1) removal of redundant rules, (2) removal of essential rules, (3) injection of explicit contradictions, and (4) application of a wide range of logical-equivalence transformations.
These transformed instances, together with the untouched base examples, form a comprehensive evaluation suite. Models such as BERT, Qwen2, and TinyLlama are fine-tuned solely on the base cases using LoRA-based parameter-efficient adaptation. They are then tested on all structural variants to assess reasoning generalisation under controlled perturbations. The evaluation module computes accuracy and deviation from the base condition ($\Delta_{\text{base}}$), enabling fine-grained analysis of model generalisation, sensitivity to missing or conflicting information, and stability under logic-preserving rewrites.
Extended Logic Rules for Multi-Step Reasoning
This chapter explores the reasoning capabilities of AI through a comprehensive case study that demonstrates the principle of “less is more” in logical deduction. We begin with a complex base example incorporating dilemma reasoning patterns, then systematically modify it in three distinct ways to reveal different aspects of AI reasoning:
  1. 1.
    Rule Reduction: Removing redundant rules while maintaining the same conclusions
  2. 2.
    Rule Equivalence: Replace multiple rules with fewer logically equivalent rules, including the laws of contraposition, implication, double negation, De Morgan, identity, and commutativity.
  3. 3.
    Rule Interference: Adding contradictory rules that disrupt the AI’s reasoning process.
All examples use consistent predicates and individuals (Anne, green, blue, cold, rough, young) to facilitate comparison across variations. The example for each variation can be found in the appendix.
Experiment
Dataset Construction To evaluate reasoning generalisation under controlled structural perturbations, we construct a synthetic dataset consisting of multi-step logical inference chains and systematically generated variants. Each base example includes a fact specifying an entity’s attribute (e.g., “Anne is green or blue”), a rule chain of five inference steps (e.g., green → cold → rough → young → nice), four questions querying intermediate and final states, and binary ground-truth answers. From each base example, we generate eleven diagnostic variants grouped into four categories:
Rule Deletion. Variant 1: removal of a logically redundant rule. Variant 2: removal of a logically essential rule. Contradictory Evidence. Variant 3: addition of an explicit contradiction in the facts. Logic-Preserving Equivalence Laws. Six single-law variants using distinct logical equivalence transformations: contrapositive, double negation, implication, De Morgan, identity, and commutativity. Multi-Law Equivalence Stacking. A final variant consisting of 2–5 equivalence-law transformations simultaneously applied to the base rule.
The dataset contains 100 base groups, of which 80 are used for training and 20 for testing; all variants are applied to all base groups, resulting in a comprehensive evaluation suite with consistent ground truth. All transformations retain natural-language surface forms to ensure that evaluation probes reasoning rather than symbolic manipulation skills.
Models Evaluated We evaluate three representative model families commonly used in reasoning research: BERT-base (devlin2018bert): a bidirectional encoder-only architecture. Qwen2-1.5B (yang2025qwen3): a decoder-only transformer representative of modern instruction-tuned LLMs. TinyLlama-1.1B (zhang2024tinyllama): an efficient LLaMA-like decoder model trained on open datasets.
Training Procedure All models are fine-tuned using LoRA (Low-Rank Adaptation) to ensure parameter-efficient training and allow direct comparison across architectures. Training is conducted on the 80 base examples only—none of the variants are seen during training. Each question is treated as an independent binary classification instance.
Input Formatting Each training instance concatenates: [facts] + [rules] + [question] with the binary label corresponding to the ground-truth answer (“T” or “F”).
Hyperparameters To maintain comparability across models, all fine-tuning uses: Batch size: 4, Learning rate: $2\times 10^{-5}$, Epochs: 10, AdamW optimizer, Maximum sequence length: 512, and LoRA rank: 8. For decoder-only models (Qwen2 and TinyLlama), the tokenizer’s pad_token is set to the eos_token to avoid padding-related errors.
Evaluation Protocol We evaluate each model on all eleven test splits: Base test set (20 unseen base examples), Rule Deletion Variants, Contradictory Evidence Variant, Six Equivalence-Law Variants, and Multi-Law Variant.
Each test contains four questions per example, producing 80 questions for the base set and 400 questions per variant. For each input, the model predicts a binary label.
During evaluation, we additionally record: The exact facts and rules presented, the equivalence laws used (if any), Question-level predictions, and a human-readable description of the structural perturbation. This allows granular error analysis of different reasoning stressors.
We use accuracy as the primary metric, reflecting the percentage of correctly answered questions in each test split. Because each question is binary and independent, accuracy provides a direct measure of reasoning correctness.
Discussion and Implications
BERT Qwen2 TinyLlama
Split Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
base 1.00 0.00 1.00 0.00 1.00 0.00
variant1 1.00 0.00 1.00 0.00 1.00 0.00
variant2 0.25 -0.75 0.25 -0.75 0.25 -0.75
variant3 0.00 -1.00 0.00 -1.00 0.00 -1.00
variant4-contrapositive 1.00 0.00 1.00 0.00 1.00 0.00
variant4-double-negation 1.00 0.00 1.00 0.00 1.00 0.00
variant4-implication 1.00 0.00 1.00 0.00 1.00 0.00
variant4-de-morgan 1.00 0.00 1.00 0.00 1.00 0.00
variant4-identity 1.00 0.00 1.00 0.00 1.00 0.00
variant4-commutativity 1.00 0.00 1.00 0.00 1.00 0.00
variant4-multi (2–5 laws) 1.00 0.00 1.00 0.00 1.00 0.00
Table 1: Accuracy and deviation from base ($\Delta$) for all models across structural variants.
Table 1 presents a unified comparison of BERT, Qwen2, and TinyLlama across all structural variants in our benchmark. Despite architectural differences, all three models exhibit an almost perfectly aligned pattern of behavior. Each model achieves 100% accuracy on the base reasoning task and retains this performance under redundant rule removal (Variant 1) as well as all logic-preserving equivalence transformations, including six single-law rewrites and multi-law (2–5) equivalence stacking. These conditions yield $\Delta=0$, confirming complete invariance to semantic-preserving logical reformulations.
In contrast, performance collapses when essential rule information is removed (Variant 2), with accuracy dropping uniformly to 25% for all models, corresponding to a $\Delta$ of 0.75-0.75. Even more strikingly, all models fail entirely under contradictory facts (Variant 3), scoring 0% accuracy and yielding the largest degradation from baseline ($\Delta=-1.00$). This cross-model consistency indicates that generalisation to logical equivalence and brittleness to structural disruptions are not tied to model size or architecture, but instead reflect systematic properties of current LLM reasoning mechanisms. The variations in our framework reveal key properties of LLM-based reasoning. By systematically manipulating the logical environment—removing rules, introducing contradictions, and applying logic-preserving equivalence transformations—we expose dimensions of generalisation and brittleness often hidden in more complex benchmarks. Together, these findings clarify what contemporary LLMs can and cannot generalise in rule-based reasoning.
Generalisation and Sensitivity
The progression from identical conclusions (Variation 1) to partially disrupted conclusions (Variation 2) to complete breakdown (Variation 3) reveals the fine-grained sensitivity of LLM reasoning to structural properties of the rule set. In Variation 1, the removal of a redundant rule leaves all models completely unaffected—indicating that LLMs can identify when a rule does not contribute new inferential content. In sharp contrast, Variation 2 demonstrates that eliminating a single essential step in the reasoning chain causes accuracy to collapse to chance levels, suggesting that LLMs cannot reconstruct missing inferential links even when the rest of the context is intact.
Variation 3 highlights a different form of fragility: all models overwhelmingly prefer rule-driven inferences over explicitly stated contradictory facts. This behavior implies that LLMs tend to follow the internal logic of the given rule system rather than adjusting their conclusions in response to new or conflicting information—an important consideration for applications involving dynamic or noisy environments.
The invariance observed across all logic-preserving rewrites—whether single-law or multi-law—shows that LLMs possess semantic stability, treating contraposition, double negation, implication re-expression, De Morgan transformations, and other equivalence laws as interchangeable. This indicates that they capture structural relationships between logically equivalent formulations, even though they falter with incomplete or inconsistent evidence. Overall, LLMs remain generalise to semantic-preserving transformations but extremely brittle to structural disruptions, a duality with important implications for AI safety: systems that appear reliable under surface-level variation may still fail when essential information is missing or contradictory, highlighting the need for methods that compensate for incomplete reasoning or reconcile conflicting evidence.
Conclusion
This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations to logical reasoning. By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolated reasoning behaviors often obscured in prior benchmarks. Across BERT, Qwen2, and LLaMA-like models, we observed perfect performance and complete invariance to semantic-preserving rewrites, but sharp failures when essential rules were removed and total collapse under explicit contradictions. These results highlight a clear gap between surface-level generalisation and deeper logical competence, underscoring the need for models that can integrate incomplete information, handle conflicting evidence, and recover missing reasoning steps. Our framework provides a foundation for improving both LLM reasoning reliability and evaluation methodology.
Appendix A Appendix
Base Example: Complex Dilemma Reasoning Structure
We begin with a comprehensive example that establishes the core reasoning pattern, drawing on a structure analogous to the Paradox of the Court [^2] The Paradox of the Court involves a contract between the teacher Protagoras and his student Euathlus, where the student only pays for lessons if he wins a court case. When Protagoras sues Euathlus for the fee, a paradox arises: if Euathlus wins, he owes nothing, but if he loses, he still avoids payment, creating a logical contradiction about the outcome., a classic logical dilemma where multiple possible paths lead to the same conclusion, despite their apparent differences.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is not young then they are not rough. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: T]
The fact “Anne is green or blue” combined with Rules 1 and 2 creates a classic dilemma: both possibilities lead to the same conclusion. This dilemma reasoning yields “Anne is cold.” Rule 3 then derives “Anne is rough” from cold, Rule 4 derives “Anne is young” from rough, and Rule 6 derives “Anne is nice” from young. Rule 5 creates a circular reinforcement but doesn’t alter the conclusions.
The logical structure can be represented as:
$(G_{a}\lor B_{a})\land(G_{a}\rightarrow C_{a})\land(B_{a}\rightarrow C_{a})\vdash C_{a}$
  • $G_{a}$: Green(Anne)
  • $B_{a}$: Blue(Anne)
  • $C_{a}$: Cold(Anne)
Variation 1: Rule Reduction with Same Conclusions
This variation demonstrates that removing redundant rules preserves the AI’s ability to reach the same conclusions, illustrating that fewer rules can be equally effective.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: T]
The reasoning proceeds identically to the base case: the dilemma from Rules 1-2 yields “Anne is cold,” Rule 3 yields rough, Rule 4 yields young, and Rule 6 yields nice. The removal of Rule 5 has no impact on the conclusions, demonstrating its redundancy.
Key Insight: AI systems that recognize this redundancy can simplify their reasoning processes without sacrificing accuracy, embodying the “less is more” principle.
The simplified logical structure becomes:
$(G_{a}\lor B_{a})\land(G_{a}\rightarrow C_{a})\land(B_{a}\rightarrow C_{a})\vdash C_{a}$
$C_{a}\land(C_{a}\rightarrow R_{a})\vdash R_{a}$
$R_{a}\land(R_{a}\rightarrow Y_{a})\vdash Y_{a}$
$Y_{a}\land(Y_{a}\rightarrow N_{a})\vdash N_{a}$
  • $G_{a}=\mathrm{Green}(Anne)$
  • $B_{a}=\mathrm{Blue}(Anne)$
  • $C_{a}=\mathrm{Cold}(Anne)$
  • $R_{a}=\mathrm{Rough}(Anne)$
  • $Y_{a}=\mathrm{Young}(Anne)$
  • $N_{a}=\mathrm{Nice}(Anne)$
Variation 2: Rule Equivalence with Different Conclusions
This variation replaces multiple rules with logically equivalent fewer rules, but interestingly leads to different conclusions due to the modified rule interactions.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule A: If someone is green or blue then they are cold. $\forall x\,((\text{Green}(x)\lor\text{Blue}(x))\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: F]
Rule A directly captures the dilemma of Rules 1 and 2, leading to “Anne is cold” with equivalent logical force. Rules 3 and 4 then proceed as before. However, without Rule 6, we cannot derive “Anne is nice,” resulting in a different conclusion for Q4.
Key Insight: While Rule A is logically equivalent to the combination of Rules 1 and 2, the overall rule set simplification changes the available inference paths, demonstrating that equivalence at the micro-level doesn’t guarantee identical macro-level conclusions.
The logical equivalence can be shown as:
$(\forall x\,(G_{x}\rightarrow C_{x})\land\forall x\,(B_{x}\rightarrow C_{x}))\equiv\forall x\,((G_{x}\lor B_{x})\rightarrow C_{x})$
  • $G_{x}=\mathrm{Green}(x)$
  • $B_{x}=\mathrm{Blue}(x)$
  • $C_{x}=\mathrm{Cold}(x)$
However, the missing Rule 6 prevents the derivation of Nice(Anne), showing that local equivalence doesn’t preserve global derivability.
Variation 3: Rule Interference with Contradictory Conclusions
This variation adds distracting and potentially contradictory rules, testing the AI’s ability to resolve conflicts and maintain coherent reasoning.
  • Facts:
    • Anne is green or blue
    • Anne is not cold or not nice
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: F]
    • Q2: Anne is rough. True/False? [Answer: F]
    • Q3: Anne is young. True/False? [Answer: F]
    • Q4: Anne is nice. True/False? [Answer: F]
Reasoning Process and Analysis
This variation introduces contradictory interference, which challenges the AI’s ability to resolve conflicts within the logical structure. The fact “Anne is green or blue,” combined with the original chain of reasoning (Rules 1-6), leads to the conclusions that Anne is cold, rough, young, and nice. However, the fact “Anne is not cold or not nice” introduces a conflict, as it implies that being nice would mean Anne is not cold. This creates an inconsistency that different AI systems might resolve differently:
  • Conservative approach: Detect inconsistency and withhold conclusions
  • Priority-based approach: Apply rule priorities or specificity heuristics
  • Paraconsistent approach: Accept some contradictions and continue reasoning
In this case, a conservative reasoning system would recognize the contradiction and potentially reject all derived conclusions, resulting in false for all questions.
The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for inconsistency detection and resolution.
The contradiction can be formally represented as:
$(C_{a}\land N_{a})\land(N_{a}\rightarrow\neg C_{a})\vdash\bot$
  • $C_{a}=\mathrm{Cold}(Anne)$
  • $N_{a}=\mathrm{Nice}(Anne)$
  • $\bot$ represents a value that is always false.
Comparative Analysis
Table 2 provides a clear comparison of how different modifications to the rule set lead to distinct conclusion patterns, highlighting the sensitivity of reasoning systems to structural changes. These variations demonstrate how even minor adjustments to the logical framework can significantly impact the reasoning process and the final outcomes.
Variations Cold Rough Young Nice
Base Example T T T T
Rule Reduction T T T T
Rule Equivalence T T T F
Rule Interference F F F F
Table 2: Comparison of conclusions across variations