### v2

# Less Is More for Multi-Step Logical Reasoning of LLM Generalisation Under Rule Removal, Paraphrasing, and Compression

Abstract
Large language models (LLMs) achieve strong performance on many natural language tasks, yet their generalisation under structured perturbations of logical rule systems remains insufficiently characterised. We present a controlled evaluation framework that probes reasoning reliability through four stress tests: (1) rule deletion, removing redundant versus essential rules from a multi-step inference chain; (2) contradictory evidence injection; (3) logic-preserving rewrites based on equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity); and (4) multi-law equivalence stacking that composes 2–5 transformations.
Across three representative model families—BERT, Qwen2, and LLaMA-like models—all models attain Acc=1.0000=1.0000 on the base split and show no degradation under redundant rule deletion. In contrast, essential rule deletion yields a pronounced decrease to near-chance performance, and injecting explicit contradictions reduces accuracy to 0.0000. Under logic-preserving rewrites, accuracy is largely preserved for single-law transformations with only small degradations in a few cases, whereas multi-law stacking exposes model-dependent sensitivity: BERT matches the base condition, TinyLlama shows only marginal degradation, and Qwen2 exhibits a substantial drop.
Overall, the results indicate that contemporary LLMs are generally stable under semantic-preserving reformulations, yet remain brittle to missing or inconsistent evidence and may degrade under composed logical transformations depending on the model family. The proposed framework provides a concise diagnostic tool for isolating these failure modes and for evaluating logical generalisation beyond surface-form variation.
Codehttps://github.com/14H034160212/lemo
Introduction
Reasoning over structured rule systems is a foundational capability for reliable decision-making and language understanding. Although large language models (LLMs) perform strongly on many benchmarks, it remains unclear to what extent such performance reflects robust logical inference as opposed to distributional heuristics. Moreover, practical reasoning settings frequently involve incomplete information, redundant premises, rephrased or logically transformed rules, and even explicit contradictions. Characterising model behaviour under these controlled perturbations is therefore important for both scientific understanding and safe deployment.
To address this gap, we introduce a controlled evaluation framework that isolates key aspects of rule-based reasoning. The benchmark is built around a canonical multi-step inference chain (e.g., green $\rightarrow$ cold $\rightarrow$ rough $\rightarrow$ young $\rightarrow$ nice) requiring multi-hop integration of rules. From each base instance, we generate targeted variants designed to probe distinct dimensions of generalisation:
Rule deletion. We remove either redundant rules (whose removal should preserve derivability) or essential rules required to complete the inference chain, thereby testing sensitivity to missing inferential links.
Contradictory evidence injection. We inject explicit facts that conflict with rule-derived conclusions, testing whether models can accommodate inconsistency rather than defaulting to a single inference path.
Logic-preserving transformations. We rewrite rules using standard equivalence laws (contraposition, double negation, implication-to-disjunction, De Morgan, identity, and commutativity). These variants preserve semantics while altering surface form, allowing us to test whether models track logical content beyond templated phrasing.
Multi-law equivalence stacking. We compose 2–5 equivalence laws simultaneously to create structurally more complex but logically equivalent rule sets, testing robustness under composed transformations.
We evaluate the framework on three model families—BERT, Qwen2, and LLaMA-like models—under identical training and evaluation pipelines. Despite architectural differences, the observed behaviour is broadly consistent:
  • Base reasoning is solved at Acc=1.0000=1.0000.
  • Redundant rule deletion yields no measurable degradation.
  • Essential rule deletion produces a marked decrease to near-chance performance.
  • Injecting explicit contradictions reduces accuracy to 0.0000.
  • Single-law equivalence rewrites are handled robustly, with only minor degradations in a small number of cases.
  • Multi-law stacking exposes model-dependent sensitivity: some models retain baseline accuracy, whereas others exhibit substantial degradation.
Together, these results highlight an asymmetry: contemporary LLMs are generally stable under semantic-preserving rewrites of rules, yet highly sensitive to missing or inconsistent evidence, and may degrade under composed logical transformations. The proposed framework provides an extensible, fine-grained diagnostic suite for isolating these failure modes, complementing existing reasoning benchmarks that often conflate multiple sources of variation.
Related Work
Research on logical reasoning in large language models (LLMs) spans multi-step deduction, logic-driven augmentation, abductive explanation, abstract reasoning assessment, generalisation analysis, and hybrid neuro-symbolic methods. A growing body of work shows that LLMs often fail even on elementary logical transformations. For example, the Reversal Curse demonstrates that models trained on “A is B” frequently fail to infer the logically equivalent “B is A,” revealing structural weaknesses in bidirectional generalisation (berglund2023reversal). Such findings motivate deeper investigation into how LLMs internalise logical relationships and under which conditions reasoning fidelity deteriorates.
Early studies like Transformers as Soft Reasoners demonstrated that transformers can perform multi-step inference over natural language rules and generalise to deeper reasoning depths (transformerClark). Subsequent work showed that this generalisation is fragile: Bao et al. found that models trained on short reasoning chains fail to extrapolate to longer or structurally shifted chains, highlighting persistent deficiencies in compositional generalisation (bao2022multi). Similar concerns are echoed in broader analyses showing that LLMs lack strong inductive biases and struggle with systematic, relational, or abstract reasoning (ijcai2024p693; Cheng2025Em; cheng2026sapp).
A parallel line of work attempts to strengthen reasoning via logic-driven augmentation. Wang et al. propose a symbolic context-extension and augmentation framework that applies equivalence laws to derive new training instances, improving model performance on ReClor and LogiQA (wang-etal-2022-logic). Bao et al. extend this idea by using Abstract Meaning Representation (AMR) to generate semantically grounded logical variants that enhance generalisation to unseen reasoning depths (bao2024). Despite their effectiveness, these methods primarily expand data distributions rather than isolating the minimal perturbations that cause reasoning failures.
Beyond deduction, abductive reasoning has been explored through frameworks such as AbductionRules, which train models to generate rules explaining surprising inputs (young-etal-2022-abductionrules). Other studies examine LLM generalisation under structured perturbations: Bao et al. show that reordering, compressing, or restructuring premises leads to substantial degradation in reasoning fidelity (baoAccess). Hybrid neuro-symbolic approaches, such as ChatLogic, integrate LLMs with Prolog-style inference mechanisms to address these limitations, although they rely heavily on external symbolic scaffolds (10650138). Complementing these evaluations, Bao’s dissertation provides a unified perspective on out-of-distribution generalisation, structural generalisation, and equivalence invariance by analysing architecture-agnostic failure modes across premise deletion, contradiction insertion, and logic-preserving rewrites (bao2025thesis).
Across these threads, prior work demonstrates meaningful progress but still lacks frameworks capable of systematically disentangling generalisation to redundant versus essential evidence, contradiction sensitivity, and invariance to different classes of logical equivalence. These unresolved limitations motivate our controlled, fine-grained evaluation framework that isolates minimal structural perturbations and reveals predictable failure modes in multi-step reasoning across LLM architectures.
Methodology
1. Synthetic Logical Environment
Base facts + rule chain generation
$\Downarrow$
2. Structural Variants
V1: Redundant rule removal
V2: Essential rule removal V3: Contradictory facts V4: Logical equivalence laws (1–5)
$\Downarrow$
3. Model Fine-Tuning & Inference
BERT / Qwen2 / TinyLlama (LoRA)
$\Downarrow$
4. Evaluation & Analysis
Accuracy, $\Delta_{\text{base}}$, variant breakdown
Figure 1: Compact system architecture of our logical reasoning generalisation framework. Each base example is transformed into structural variations (rule deletion, contradiction injection, and logic-equivalence rewrites). Models are trained on base data only and evaluated across all variants to measure generalisation and sensitivity.
The system architecture in Figure 1 outlines the full pipeline used to construct, transform, and evaluate logical reasoning tasks. The process begins with the generation of synthetic logical environments containing base facts and a multi-step rule chain designed to support deterministic deductive inferences. From each base environment, we systematically derive multiple structural variants that probe different dimensions of reasoning generalisation: (1) removal of redundant rules, (2) removal of essential rules, (3) injection of explicit contradictions, and (4) application of a wide range of logical-equivalence transformations.
These transformed instances, together with the untouched base examples, form a comprehensive evaluation suite. Models such as BERT, Qwen2, and TinyLlama are fine-tuned solely on the base cases using LoRA-based parameter-efficient adaptation. They are then tested on all structural variants to assess reasoning generalisation under controlled perturbations. The evaluation module computes accuracy and deviation from the base condition ($\Delta_{\text{base}}$), enabling fine-grained analysis of model generalisation, sensitivity to missing or conflicting information, and stability under logic-preserving rewrites.
Extended Logic Rules for Multi-Step Reasoning
This chapter explores the reasoning capabilities of AI through a comprehensive case study that demonstrates the principle of “less is more” in logical deduction. We begin with a complex base example incorporating dilemma reasoning patterns, then systematically modify it in three distinct ways to reveal different aspects of AI reasoning:
  1. 1.
    Rule Reduction: Removing redundant rules while maintaining the same conclusions
  2. 2.
    Rule Equivalence: Replace multiple rules with fewer logically equivalent rules, including the laws of contraposition, implication, double negation, De Morgan, identity, and commutativity.
  3. 3.
    Rule Interference: Adding contradictory rules that disrupt the AI’s reasoning process.
All examples use consistent predicates and individuals (Anne, green, blue, cold, rough, young) to facilitate comparison across variations. The example for each variation can be found in the appendix.
Experiment
Dataset Construction To evaluate reasoning generalisation under controlled structural perturbations, we construct a synthetic dataset consisting of multi-step logical inference chains and systematically generated variants. Each base example includes a fact specifying an entity’s attribute (e.g., “Anne is green or blue”), a rule chain of five inference steps (e.g., green → cold → rough → young → nice), four questions querying intermediate and final states, and binary ground-truth answers. From each base example, we generate eleven diagnostic variants grouped into four categories:
Rule Deletion. Variant 1: removal of a logically redundant rule. Variant 2: removal of a logically essential rule. Contradictory Evidence. Variant 3: addition of an explicit contradiction in the facts. Logic-Preserving Equivalence Laws. Six single-law variants using distinct logical equivalence transformations: contrapositive, double negation, implication, De Morgan, identity, and commutativity. Multi-Law Equivalence Stacking. A final variant consisting of 2–5 equivalence-law transformations simultaneously applied to the base rule.
The dataset contains 100 base groups, of which 80 are used for training and 20 for testing; all variants are applied to all base groups, resulting in a comprehensive evaluation suite with consistent ground truth. All transformations retain natural-language surface forms to ensure that evaluation probes reasoning rather than symbolic manipulation skills.
Models Evaluated We evaluate three representative model families commonly used in reasoning research: BERT-base (devlin2018bert): a bidirectional encoder-only architecture. Qwen2-1.5B (yang2025qwen3): a decoder-only transformer representative of modern instruction-tuned LLMs. TinyLlama-1.1B (zhang2024tinyllama): an efficient LLaMA-like decoder model trained on open datasets.
Training Procedure All models are fine-tuned using LoRA (Low-Rank Adaptation) to ensure parameter-efficient training and allow direct comparison across architectures. Training is conducted on the 80 base examples only—none of the variants are seen during training. Each question is treated as an independent binary classification instance.
Input Formatting Each training instance concatenates: [facts] + [rules] + [question] with the binary label corresponding to the ground-truth answer (“T” or “F”).
Hyperparameters To maintain comparability across models, all fine-tuning uses: Batch size: 4, Learning rate: $2\times 10^{-5}$, Epochs: 10, AdamW optimizer, Maximum sequence length: 512, and LoRA rank: 8. For decoder-only models (Qwen2 and TinyLlama), the tokenizer’s pad_token is set to the eos_token to avoid padding-related errors.
Evaluation Protocol We evaluate each model on all eleven test splits: Base test set (20 unseen base examples), Rule Deletion Variants, Contradictory Evidence Variant, Six Equivalence-Law Variants, and Multi-Law Variant.
Each test contains four questions per example, producing 80 questions for the base set and 400 questions per variant. For each input, the model predicts a binary label.
During evaluation, we additionally record: The exact facts and rules presented, the equivalence laws used (if any), Question-level predictions, and a human-readable description of the structural perturbation. This allows granular error analysis of different reasoning stressors.
We use accuracy as the primary metric, reflecting the percentage of correctly answered questions in each test split. Because each question is binary and independent, accuracy provides a direct measure of reasoning correctness.
Discussion and Implications
BERT Qwen2 TinyLlama
Split Acc $\Delta$ Acc $\Delta$ Acc $\Delta$
base 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant1 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant2 0.2950 -0.7050 0.2500 -0.7500 0.2500 -0.7500
variant3 0.0000 -1.0000 0.0000 -1.0000 0.0000 -1.0000
variant4-contrapositive 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-double-negation 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-implication 1.0000 0.0000 0.9525 -0.0475 1.0000 0.0000
variant4-de-morgan 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-identity 1.0000 0.0000 1.0000 0.0000 1.0000 0.0000
variant4-commutativity 0.9925 -0.0075 1.0000 0.0000 1.0000 0.0000
variant4-multi (2–5 laws) 1.0000 0.0000 0.6450 -0.3550 0.9925 -0.0075
Table 1: Accuracy and deviation from base ($\Delta$) for all models across structural variants.
Table 1 reports accuracy (Acc) and deviation from the base condition ($\Delta=\mathrm{Acc}_{\mathrm{variant}}-\mathrm{Acc}_{\mathrm{base}}$) for BERT, Qwen2, and TinyLlama across all structural variants. All models achieve Acc=1.0000=1.0000 on the base split and exhibit no degradation under redundant rule removal (Variant 1; $\Delta=0$). By contrast, removing an essential rule (Variant 2) yields a substantial drop (BERT: 0.2950, $\Delta=-0.7050$; Qwen2/TinyLlama: 0.2500, $\Delta=-0.7500$), indicating strong sensitivity to missing inferential links. Injecting explicit contradictory facts (Variant 3) reduces accuracy to 0.0000 for all models ($\Delta=-1.0000$), suggesting that the models do not reliably revise conclusions in the presence of inconsistency.
For logic-preserving single-law rewrites (Variant 4), performance is largely preserved: most conditions remain at 1.0000 accuracy, with only small degradations in a few settings (e.g., BERT under commutativity: 0.9925, $\Delta=-0.0075$; Qwen2 under implication-to-disjunction: 0.9525, $\Delta=-0.0475$). Under multi-law equivalence stacking (2–5 laws), robustness becomes model-dependent: BERT matches the base condition (1.0000), TinyLlama shows only a marginal reduction (0.9925, $\Delta=-0.0075$), whereas Qwen2 exhibits a pronounced decrease (0.6450, $\Delta=-0.3550$). This contrast suggests that compositional logical rewrites can introduce effective difficulty beyond single transformations, even when semantic equivalence is preserved.
Generalisation and Sensitivity
Variants 1–3 isolate model sensitivity to perturbations that preserve versus disrupt derivability. In Variant 1 (redundant rule deletion), all models match the base accuracy, consistent with the expectation that removing unused rules should not affect entailment. In Variant 2 (essential rule deletion), accuracy drops to near-chance levels, suggesting limited premise tracking: models often behave as if the missing link were still present and tend to “complete” the original inference pattern even when intermediate support is removed. This behaviour is consistent with reliance on learned chain templates rather than verifying support for each queried conclusion. In Variant 3 (contradictory evidence), all models score 0.0000, indicating that they do not reliably adopt contradiction-aware conservative reasoning under the benchmark’s scoring semantics.
For logic-preserving rewrites (Variant 4), performance is largely preserved for single-law transformations, with only minor degradations in a few cases (e.g., implication-to-disjunction or commutativity). However, multi-law stacking exposes model-dependent robustness: BERT remains at the base level, TinyLlama shows only marginal reduction, whereas Qwen2 degrades substantially. This suggests that composing multiple equivalence transformations introduces an effective distribution shift beyond isolated rewrites, likely through increased operator density and structural complexity despite semantic equivalence.
Implications and Diagnostic Value
The variant-wise profile provides a compact diagnostic of three failure modes: (i) missing-link sensitivity under essential rule deletion, indicating weak verification of minimal supporting premises; (ii) inconsistency handling under contradiction injection, where models fail to revise or withhold conclusions when counter-evidence is present; and (iii) compositional sensitivity, where multi-law rewrites can degrade performance even when each rewrite is semantics-preserving. The strong gap between single-law and multi-law conditions cautions against equating robustness to isolated perturbations with robustness under composition. These diagnostics motivate targeted interventions such as deletion-aware training, explicit proof/rationale verification, and stress-testing robustness under composed transformations rather than single perturbations alone.
Conclusion
This work introduced a controlled framework for evaluating LLM generalisation under systematic perturbations of rule-based reasoning. By constructing multi-step inference chains and generating variants involving rule deletion, contradiction injection, and multiple families of logic-preserving transformations, we isolate behaviours that are often obscured in more heterogeneous benchmarks. Across BERT, Qwen2, and LLaMA-like models, we observe Acc=1.0000=1.0000 on the base split and no degradation under redundant rule deletion, but pronounced failures under essential rule deletion and uniform breakdown under explicit contradictions. For equivalence-based rewrites, performance is largely preserved for single-law transformations, while composed (multi-law) rewrites reveal model-dependent sensitivity. These findings underscore the gap between robustness to semantic-preserving reformulations and reliable reasoning under missing, inconsistent, or compositionally transformed evidence, and motivate methods that explicitly address premise tracking and inconsistency management.
Appendix A Appendix
Base Example: Complex Dilemma Reasoning Structure
We begin with a comprehensive example that establishes the core reasoning pattern, drawing on a structure analogous to the Paradox of the Court [^2] The Paradox of the Court involves a contract between the teacher Protagoras and his student Euathlus, where the student only pays for lessons if he wins a court case. When Protagoras sues Euathlus for the fee, a paradox arises: if Euathlus wins, he owes nothing, but if he loses, he still avoids payment, creating a logical contradiction about the outcome., a classic logical dilemma where multiple possible paths lead to the same conclusion, despite their apparent differences.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: T]
The fact “Anne is green or blue” combined with Rules 1 and 2 creates a classic dilemma: both possibilities lead to the same conclusion. This dilemma reasoning yields “Anne is cold.” Rule 3 then derives “Anne is rough” from cold, Rule 4 derives “Anne is young” from rough, and Rule 6 derives “Anne is nice” from young. Rule 5 creates a circular reinforcement but doesn’t alter the conclusions.
The logical structure can be represented as:
$(G_{a}\lor B_{a})\land(G_{a}\rightarrow C_{a})\land(B_{a}\rightarrow C_{a})\vdash C_{a}$
  • $G_{a}$: Green(Anne)
  • $B_{a}$: Blue(Anne)
  • $C_{a}$: Cold(Anne)
Variation 1: Rule Reduction with Same Conclusions
This variation demonstrates that removing redundant rules preserves the AI’s ability to reach the same conclusions, illustrating that fewer rules can be equally effective.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: T]
The reasoning proceeds identically to the base case: the dilemma from Rules 1-2 yields “Anne is cold,” Rule 3 yields rough, Rule 4 yields young, and Rule 6 yields nice. The removal of Rule 5 has no impact on the conclusions, demonstrating its redundancy.
Key Insight: AI systems that recognize this redundancy can simplify their reasoning processes without sacrificing accuracy, embodying the “less is more” principle.
The simplified logical structure becomes:
$(G_{a}\lor B_{a})\land(G_{a}\rightarrow C_{a})\land(B_{a}\rightarrow C_{a})\vdash C_{a}$
$C_{a}\land(C_{a}\rightarrow R_{a})\vdash R_{a}$
$R_{a}\land(R_{a}\rightarrow Y_{a})\vdash Y_{a}$
$Y_{a}\land(Y_{a}\rightarrow N_{a})\vdash N_{a}$
  • $G_{a}=\mathrm{Green}(Anne)$
  • $B_{a}=\mathrm{Blue}(Anne)$
  • $C_{a}=\mathrm{Cold}(Anne)$
  • $R_{a}=\mathrm{Rough}(Anne)$
  • $Y_{a}=\mathrm{Young}(Anne)$
  • $N_{a}=\mathrm{Nice}(Anne)$
Variation 2: Rule Equivalence with Different Conclusions
This variation replaces multiple rules with logically equivalent fewer rules, but interestingly leads to different conclusions due to the modified rule interactions.
  • Facts:
    • Anne is green or blue
  • Rules:
    • Rule A: If someone is green or blue then they are cold. $\forall x\,((\text{Green}(x)\lor\text{Blue}(x))\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: T]
    • Q2: Anne is rough. True/False? [Answer: T]
    • Q3: Anne is young. True/False? [Answer: T]
    • Q4: Anne is nice. True/False? [Answer: F]
Rule A directly captures the dilemma of Rules 1 and 2, leading to “Anne is cold” with equivalent logical force. Rules 3 and 4 then proceed as before. However, without Rule 6, we cannot derive “Anne is nice,” resulting in a different conclusion for Q4.
Key Insight: While Rule A is logically equivalent to the combination of Rules 1 and 2, the overall rule set simplification changes the available inference paths, demonstrating that equivalence at the micro-level doesn’t guarantee identical macro-level conclusions.
The logical equivalence can be shown as:
$(\forall x\,(G_{x}\rightarrow C_{x})\land\forall x\,(B_{x}\rightarrow C_{x}))\equiv\forall x\,((G_{x}\lor B_{x})\rightarrow C_{x})$
  • $G_{x}=\mathrm{Green}(x)$
  • $B_{x}=\mathrm{Blue}(x)$
  • $C_{x}=\mathrm{Cold}(x)$
However, the missing Rule 6 prevents the derivation of Nice(Anne), showing that local equivalence doesn’t preserve global derivability.
Variation 3: Rule Interference with Contradictory Conclusions
This variation adds distracting and potentially contradictory rules, testing the AI’s ability to resolve conflicts and maintain coherent reasoning.
  • Facts:
    • Anne is green or blue
    • Anne is not cold or not nice
  • Rules:
    • Rule 1: If someone is green then they are cold. $\forall x\,(\text{Green}(x)\rightarrow\text{Cold}(x))$
    • Rule 2: If someone is blue then they are cold. $\forall x\,(\text{Blue}(x)\rightarrow\text{Cold}(x))$
    • Rule 3: If someone is cold then they are rough. $\forall x\,(\text{Cold}(x)\rightarrow\text{Rough}(x))$
    • Rule 4: If someone is rough then they are young. $\forall x\,(\text{Rough}(x)\rightarrow\text{Young}(x))$
    • Rule 5: If someone is young then they are cold. $\forall x\,(\text{Young}(x)\rightarrow\text{Cold}(x))$
    • Rule 6: If someone is young then they are nice. $\forall x\,(\text{Young}(x)\rightarrow\text{Nice}(x))$
  • Questions:
    • Q1: Anne is cold. True/False? [Answer: F]
    • Q2: Anne is rough. True/False? [Answer: F]
    • Q3: Anne is young. True/False? [Answer: F]
    • Q4: Anne is nice. True/False? [Answer: F]
Reasoning Process and Analysis
This variation introduces contradictory interference, which challenges the AI’s ability to resolve conflicts within the logical structure. The fact “Anne is green or blue,” combined with the original chain of reasoning (Rules 1-6), leads to the conclusions that Anne is cold, rough, young, and nice. However, the fact “Anne is not cold or not nice” introduces a conflict, as it implies that being nice would mean Anne is not cold. This creates an inconsistency that different AI systems might resolve differently:
  • Conservative approach: Detect inconsistency and withhold conclusions
  • Priority-based approach: Apply rule priorities or specificity heuristics
  • Paraconsistent approach: Accept some contradictions and continue reasoning
In this case, a conservative reasoning system would recognize the contradiction and potentially reject all derived conclusions, resulting in false for all questions.
The addition of interfering rules not only tests the AI’s ability to ignore distractions but also its capacity for inconsistency detection and resolution.
The contradiction can be formally represented as:
$(C_{a}\land N_{a})\land(N_{a}\rightarrow\neg C_{a})\vdash\bot$
  • $C_{a}=\mathrm{Cold}(Anne)$
  • $N_{a}=\mathrm{Nice}(Anne)$
  • $\bot$ represents a value that is always false.
Comparative Analysis
Table 2 provides a clear comparison of how different modifications to the rule set lead to distinct conclusion patterns, highlighting the sensitivity of reasoning systems to structural changes. These variations demonstrate how even minor adjustments to the logical framework can significantly impact the reasoning process and the final outcomes.
Variations Cold Rough Young Nice
Base Example T T T T
Rule Reduction T T T T
Rule Equivalence T T T F
Rule Interference F F F F
Table 2: Comparison of conclusions across variations